Jane hiked 3 5/8 miles. Diana kiked 1 3/4 miles. How much farther did Jane hike than Diana?

Found 2 solutions by rfer, checkley79:
Answer by rfer(16322)

29/8-7/4=
29/8-14/8=15/8=1 7/8

Answer by checkley79(3341)

3 5/8-1 3/4
(3*8+5)/8-(1*4+3)/4
(24+5)/8-(4+3)/4
29/8-7/4
(29-7*2)/8
(29-14)/8
15/8=1 7/8 ANS.